OpenStudy (anonymous):

Given tan (theta)= -2 and (pi/ 2) < theta < pi; find cos 2(theta)

OpenStudy (loser66):

@Mertsj

OpenStudy (anonymous):



OpenStudy (anonymous):

find the hypotenuse via pythagoras, or in your head, get \(\sqrt{2^2+1^2}=\sqrt{5}\)

OpenStudy (anonymous):

that tells you \[\cos(\theta)=-\frac{1}{\sqrt{5}}\]

OpenStudy (anonymous):

then use the formula \[\cos(2\theta)=2\cos^2(\theta)-1\] to finish the problem

OpenStudy (loser66):

thanks @satellite73

OpenStudy (loser66):

@edualvarez14 if you don't know why cos = -(.....) , just speak out, ok? and it needs some more steps to get the final answer, do if you can.